# Reveal Hidden Pitfalls and Navigate Next Generation of Vector Similarity Search from Task-Centric Views

### v1

@@ 3. Resource Collection for Funnel Analysis
Image
Classification
ImageNet-DINOv2 1,281,167 768 DINOv2 (Oquab et al., 2024) Image Class Label Recall@K
@@
Face
Recognition
Glint360K-ViT 17,091,649 512 ArcFace-IR101 (Deng et al., 2019) Identity ID Label Recall@K
@@
Text
Retrieval
BookCorpus 9,250,529 1024 Stella_en_1.5B_v5 Paragraph ID Hit@K
Recommendation
Commerce 99,085,171 48 ResFlow (Fu et al., 2024) Item ID Matching Score@K
@@ 3.1. Datasets for Important Application
Data Collection. We adopt the open-source Glint360K dataset (gli, 2023) as the representative benchmark for face recognition. This large-scale dataset contains over 17 million images spanning 360,232 identities, curated from public sources and annotated via a combination of manual and automated methods. For query construction in $\mathsf{Iceberg}$, we randomly sample one face image per identity, with the first 20,000 identities used as the test set.
@@
Inner Product
Fargo Partition-based $O(nd)$ $O(\beta nd)$, $\beta\ll 1$ Low.
@@
Euclidean Distance
RaBitQ Partition-based $O(2(nd^{2}+nd))$ $O(c_{1}d\log\log d)$ High.
@@ 4.2. Impact of the Information Loss Funnel
Metric Method
ImageNet-DINOv2
(Label Recall@100)
Glink360K-ViT
(Label Recall@100)
IP
Fargo 71.38% 90.3%
@@
O1 (Embedding Process): Information loss first occurs during the transformation from raw data to embeddings. This loss can untie the semantic similarity with spatial closeness under given distance metrics, making the pursue for high synthetic recall meaningless. Our experiments reveal a consistent gap between synthetic metrics and true downstream task performance. First, as shown in Table 4, we observe that downstream utility (label recall@100) exhibits a natural upper bound, even when synthetic recall approaches 100%. For instance, on the ImageNet dataset with DINOv2 embeddings, label recall@100 saturates at 71% despite vector recall@100 reaching 99.9%. A similar bottleneck is observed in the Glink360K-ViT dataset, where label recall@100 plateaus at 91% while synthetic recall is nearly perfect. This indicates that high $\mathsf{VSS}$ synthetic accuracy does not necessarily translate into better task performance. Second, on the Commerce dataset, we find a non-monotonic relationship between synthetic metrics accuracy and downstream task-centric utility (Figure 5). As synthetic recall improves, the matching score@100—a task-specific measure of user interest—initially increases but then declines. This counterintuitive trend shows that relying too heavily on synthetic measures without considering their practical implications can lead to misleading conclusions. Finally, Table 5 shows that downstream task-centric utility is highly sensitive to the choice of embedding model. Using DINOv2, EVA02, ConvNeXt, and AlexNet on ImageNet, we observe significant differences in label recall@100 under the same retrieval setting (all with 99% synthetic recall). While DINOv2 achieves only 71%, EVA02 and ConvNeXt reach 85% and 84%, respectively. Furthermore, representations from the old embedding model diverge more from task-centric retrieval targets. AlexNet-based embeddings achieved only about 21% accuracy on label recall, suggesting that neglecting downstream evaluation on outdated embeddings leads to increasingly misaligned and less meaningful retrieval results.
@@
Figure 6 reports results across four datasets: ImageNet-EVA02, ImgeNet-ConvNeXt, BookCorpus, and Glink360-IR101, demonstrating how the similarity metric choice radically impacts downstream task-centric performance. On ImageNet, with ConvNeXt embeddings, $\mathsf{ANNS}$ using Euclidean distance scores roughly 4% higher label recall@100 than $\mathsf{MIPS}$, even though both are run at synthetic recall@100 around 99.9%. The problem is more pronounced with the EVA02 model: even though $\mathsf{MIPS}$ achieves nearly perfect synthetic recall@100 (99.99%), actual task-centric label recall@100 is stubbornly below 1%, reflecting a near-complete semantic mismatch. Analogously, on the text retrieval dataset BookCorpus, $\mathsf{MIPS}$ manages less than 50% task-centric hit recall, whereas Euclidean-based $\mathsf{ANNS}$ attains 100%. However, the Glink360-IR101 dataset shows the opposite. Here, $\mathsf{MIPS}$ comfortably outperforms Euclidean-based $\mathsf{ANNS}$ on task-centric metrics: with Glink360K-IR101 embeddings, $\mathsf{MIPS}$ performs 98% label recall@100, whereas HNSW ($\mathsf{ANNS}$ methods) tops out at 92%. Given the high-stakes nature of applications such as identity verification, such performance discrepancies result in critical system failures—even with high synthetic recall.
These differences in performance are due to mismatched similarity measures and embedding space geometries, controlled through loss design and metric affinity. Models such as ConvNeXt are trained with cross-entropy loss, which promotes embeddings within the same class to have tight clusters in Euclidean space. Euclidean distance naturally captures this structure. By contrast, $\mathsf{MIPS}$ is sensitive to both direction and norm of vectors and does not respect this clustering, retrieving semantically unrelated neighbors despite high synthetic recall. Glink360K-IR101 embeddings are trained with ArcFace, which employs an Additive Angular Margin Loss to foster a highly discriminative angular structure—small intra-class angles and large inter-class angles. This structure is naturally captured by inner product similarity, and thus $\mathsf{MIPS}$ is the correct choice and we see why it has superior task-centric performance on Glink360. Combined, these results highlight an important takeaway: Evaluating synthetic recall with a single similarity metric (like Euclidean distance) can be misleading and, unless that metric is aligned with the training loss, largely meaningless. Metric choice materially affects outcomes: metric misuse not only incurs drastic information loss but also can make high-synthetic recall $\mathsf{VSS}$ systems functionally useless in real-world downstream tasks.
@@
O-3.2: Wrong choice leads to performance degradation. Beyond the query speed disparities on synthetic metric, the impact of selecting an unsuitable method is more critical—it can constrain synthetic metric recall and downstream task-centric performance, ultimately causing significant information loss. This phenomenon is particularly pronounced in the domain of $\mathsf{MIPS}$. Due to the lack of rigorous theoretical foundations, some methods suffer from synthetic metric recall bottlenecks and performance degradation on certain datasets. For example, while the Mobius algorithm performs well on some datasets (e.g., BookCorpus), it fails to exceed 50% accuracy on synthetic metric recall@100 on the Glink360K-IR101 dataset. Similarly, partition-based methods such as $\mathsf{ScaNN}$ exhibit noticeable performance drops on the Commerce dataset when the target synthetic metric recall exceeds 90%. These findings highlight that variations in data distribution can lead to different degrees of information loss across methods.
@@ 4.5. Insights & Future Directions
O9 (Distribution-aware Improvement) The heterogeneous pattern distributions of diverse vector datasets inherently challenge traditional $\mathsf{VSS}$ methods that rely on fixed construction strategies. Such methods have poor generalizability to datasets with varied geometric properties. For example, in the Glink360K dataset, as shown in Figure 9(left), K-means clustering is performed in the original space, followed by PCA methods for 3D visualization, where data points are separated by cluster labels. The dataset showcases extreme angular separation: inter-class distances are large, while intra-class distances have high variability, thus severely crippling the usefulness of graph-based indices. For the BookCorpus dataset, an analogous dimensionality reduction method was used, with the data points colored according to their norms. Figure 9(right) shows that the average angle between vectors is kept within 60 degrees, and the data is scattered along multiple norm directions. In such scenarios, graph-based methods can effectively leverage the triangle inequality for pruning edges. However, partition-based approaches struggle to create clear boundaries between partitions, making it difficult to exactly identify border points. This limitation negatively impacts their ability to achieve high synthetic metric recall efficiently.
@@
Figure 9. Visualization of data distributions: Left – Glink360K-ViT dataset; Right – BookCorpus dataset.

### v2

@@ 3. Resource Collection for Funnel Analysis
Image Classification ImageNet-DINOv2 1,281,167 768 DINOv2 (Oquab et al., 2024) Image Class Label Recall@K
@@
Face Recognition Glint360K-ViT 17,091,649 512 ArcFace-IR101 (Deng et al., 2019) Identity ID Label Recall@K
@@
Text Retrieval BookCorpus 9,250,529 1024 Stella_en_1.5B_v5 Paragraph ID Hit@K
Recommendation Commerce 99,085,171 48 ResFlow (Fu et al., 2024) Item ID Matching Score@K
@@ 3.1. Datasets for Important Application
Data Collection. We adopt the open-source Glint360K dataset (An et al., 2021) as the representative benchmark for face recognition. This large-scale dataset contains over 17 million images spanning 360,232 identities, curated from public sources and annotated via a combination of manual and automated methods. For query construction in $\mathsf{Iceberg}$, we randomly sample one face image per identity, with the first 20,000 identities used as the test set.
@@
Inner Product Fargo Partition-based $O(nd)$ $O(\beta nd)$, $\beta\ll 1$ Low.
@@
Euclidean Distance RaBitQ Partition-based $O(2(nd^{2}+nd))$ $O(c_{1}d\log\log d)$ High.
@@ 4.2. Impact of the Information Loss Funnel
Metric Method ImageNet-DINOv2 (Label Recall@100) Glint360K-ViT (Label Recall@100)
IP Fargo 71.38% 90.3%
@@
O1 (Embedding Process): Information loss first occurs during the transformation from raw data to embeddings. This loss can untie the semantic similarity with spatial closeness under given distance metrics, making the pursue for high synthetic recall meaningless. Our experiments reveal a consistent gap between synthetic metrics and true downstream task performance. First, as shown in Table 4, we observe that downstream utility (label recall@100) exhibits a natural upper bound, even when synthetic recall approaches 100%. For instance, on the ImageNet dataset with DINOv2 embeddings, label recall@100 saturates at 71% despite vector recall@100 reaching 99.9%. A similar bottleneck is observed in the Glint360K-ViT dataset, where label recall@100 plateaus at 91% while synthetic recall is nearly perfect. This indicates that high $\mathsf{VSS}$ synthetic accuracy does not necessarily translate into better task performance. Second, on the Commerce dataset, we find a non-monotonic relationship between synthetic metrics accuracy and downstream task-centric utility (Figure 5). As synthetic recall improves, the matching score@100—a task-specific measure of user interest—initially increases but then declines. This counterintuitive trend shows that relying too heavily on synthetic measures without considering their practical implications can lead to misleading conclusions. Finally, Table 5 shows that downstream task-centric utility is highly sensitive to the choice of embedding model. Using DINOv2, EVA02, ConvNeXt, and AlexNet on ImageNet, we observe significant differences in label recall@100 under the same retrieval setting (all with 99% synthetic recall). While DINOv2 achieves only 71%, EVA02 and ConvNeXt reach 85% and 84%, respectively. Furthermore, representations from the old embedding model diverge more from task-centric retrieval targets. AlexNet-based embeddings achieved only about 21% accuracy on label recall, suggesting that neglecting downstream evaluation on outdated embeddings leads to increasingly misaligned and less meaningful retrieval results.
@@
Figure 6 reports results across four datasets: ImageNet-EVA02, ImgeNet-ConvNeXt, BookCorpus, and Glint360-IR101, demonstrating how the similarity metric choice radically impacts downstream task-centric performance. On ImageNet, with ConvNeXt embeddings, $\mathsf{ANNS}$ using Euclidean distance scores roughly 4% higher label recall@100 than $\mathsf{MIPS}$, even though both are run at synthetic recall@100 around 99.9%. The problem is more pronounced with the EVA02 model: even though $\mathsf{MIPS}$ achieves nearly perfect synthetic recall@100 (99.99%), actual task-centric label recall@100 is stubbornly below 1%, reflecting a near-complete semantic mismatch. Analogously, on the text retrieval dataset BookCorpus, $\mathsf{MIPS}$ manages less than 50% task-centric hit recall, whereas Euclidean-based $\mathsf{ANNS}$ attains 100%. However, the Glint360-IR101 dataset shows the opposite. Here, $\mathsf{MIPS}$ comfortably outperforms Euclidean-based $\mathsf{ANNS}$ on task-centric metrics: with Glint360K-IR101 embeddings, $\mathsf{MIPS}$ performs 98% label recall@100, whereas HNSW ($\mathsf{ANNS}$ methods) tops out at 92%. Given the high-stakes nature of applications such as identity verification, such performance discrepancies result in critical system failures—even with high synthetic recall.
These differences in performance are due to mismatched similarity measures and embedding space geometries, controlled through loss design and metric affinity. Models such as ConvNeXt are trained with cross-entropy loss, which promotes embeddings within the same class to have tight clusters in Euclidean space. Euclidean distance naturally captures this structure. By contrast, $\mathsf{MIPS}$ is sensitive to both direction and norm of vectors and does not respect this clustering, retrieving semantically unrelated neighbors despite high synthetic recall. Glint360K-IR101 embeddings are trained with ArcFace, which employs an Additive Angular Margin Loss to foster a highly discriminative angular structure—small intra-class angles and large inter-class angles. This structure is naturally captured by inner product similarity, and thus $\mathsf{MIPS}$ is the correct choice and we see why it has superior task-centric performance on Glint360. Combined, these results highlight an important takeaway: Evaluating synthetic recall with a single similarity metric (like Euclidean distance) can be misleading and, unless that metric is aligned with the training loss, largely meaningless. Metric choice materially affects outcomes: metric misuse not only incurs drastic information loss but also can make high-synthetic recall $\mathsf{VSS}$ systems functionally useless in real-world downstream tasks.
@@
O-3.2: Wrong choice leads to performance degradation. Beyond the query speed disparities on synthetic metric, the impact of selecting an unsuitable method is more critical—it can constrain synthetic metric recall and downstream task-centric performance, ultimately causing significant information loss. This phenomenon is particularly pronounced in the domain of $\mathsf{MIPS}$. Due to the lack of rigorous theoretical foundations, some methods suffer from synthetic metric recall bottlenecks and performance degradation on certain datasets. For example, while the Mobius algorithm performs well on some datasets (e.g., BookCorpus), it fails to exceed 50% accuracy on synthetic metric recall@100 on the Glint360K-IR101 dataset. Similarly, partition-based methods such as $\mathsf{ScaNN}$ exhibit noticeable performance drops on the Commerce dataset when the target synthetic metric recall exceeds 90%. These findings highlight that variations in data distribution can lead to different degrees of information loss across methods.
@@ 4.5. Insights & Future Directions
O9 (Distribution-aware Improvement) The heterogeneous pattern distributions of diverse vector datasets inherently challenge traditional $\mathsf{VSS}$ methods that rely on fixed construction strategies. Such methods have poor generalizability to datasets with varied geometric properties. For example, in the Glint360K dataset, as shown in Figure 9(left), K-means clustering is performed in the original space, followed by PCA methods for 3D visualization, where data points are separated by cluster labels. The dataset showcases extreme angular separation: inter-class distances are large, while intra-class distances have high variability, thus severely crippling the usefulness of graph-based indices. For the BookCorpus dataset, an analogous dimensionality reduction method was used, with the data points colored according to their norms. Figure 9(right) shows that the average angle between vectors is kept within 60 degrees, and the data is scattered along multiple norm directions. In such scenarios, graph-based methods can effectively leverage the triangle inequality for pruning edges. However, partition-based approaches struggle to create clear boundaries between partitions, making it difficult to exactly identify border points. This limitation negatively impacts their ability to achieve high synthetic metric recall efficiently.
@@
Figure 9. Visualization of data distributions: Left – Glint360K-ViT dataset; Right – BookCorpus dataset.